OpenStudy (anonymous):

PLEASE HELP ASAP! If a>1 , then which of the following statements is always true?

OpenStudy (anonymous):

A. \[\sqrt[4]{a} \ > \ \sqrt[3]{a}\] B. \[a \ > \ \sqrt[3]{a}\] C. \[\sqrt[3]{a} \ < \ 0\] D. \[a \ < \ \sqrt[3]{a}\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi Can someone please help me :(

OpenStudy (anonymous):

@dpaInc